OpenStudy (anonymous):

How many ways can Hannah give 3 of her 8 T-shirts to her younger sister? A. 40,314 B. 120 C. 336 D. 56

OpenStudy (raden):

use the 8C3

OpenStudy (raden):

now, can u evaluate the value of 8C3 ?

OpenStudy (raden):

hint : nCr = n!/r!(n-r)!

OpenStudy (anonymous):

hmmmmm, I still am confused.

OpenStudy (anonymous):

what's the n, C, and r?

OpenStudy (raden):

actually, it is the general formula of combination. also u have know about the "factorial", like 3! = 3 * 2 * 1 = 6

OpenStudy (anonymous):

ok. Can u help me solve the variables? what they should be? I was never good with this.

OpenStudy (raden):

8C3, in this case knowed n=8 and r=3. by using the formula above, it can be 8C3 = 8!/3!(8-3)! = 8!/3!5! = 8 * 7 * 6 * 5! / 3*2*1*5! (cancel out the 5!) = 8 * 7 * 6 / 3 * 2 * 1 = 8*7*6/6 (cancel out the 6) = 8 * 7 = 56

OpenStudy (raden):

so, there are 56 ways

OpenStudy (anonymous):

thank you for your help :)